# On the torsion growth in quadratic number fields for elliptic curves defined over the rationals

### v1

@@ Abstract.
This work was supported by the Ministerio de Ciencia e Innovación under Project PID2020-114613GB-I00 (MCIN/AEI/10.13039/501100011033).
@@ 1. Introduction: The problem
This paper deals with properties of the torsion subgroup of elliptic curves defined over the rationals under quadratic field extensions, a topic that was first studied in [2, 3, 9] and has been extensively researched and generalized to higher extensions since.
@@
  • Whenever we consider a quadratic number field written as $K=\mathbb{Q}(\sqrt{d})$ we will assume $d$ is square-free.
Let $E$ be an elliptic curve defined over $\mathbb{Q}$. Let $\ell$ be a prime number and let us write:
@@
As it is well–known [10], the action of the absolute Galois group $\mathrm{G}_{\mathbb{Q}}=\mathop{\mathrm{Gal}}\nolimits(\overline{\mathbb{Q}}/\mathbb{Q})$ on $E[\ell]$ defines a mod $\ell$ Galois representation
The extension $\mathbb{Q}(E[\ell])/\mathbb{Q}$ is Galois, with $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[\ell])/\mathbb{Q})\simeq\overline{\rho}_{E,\ell}(\mathrm{G}_{\mathbb{Q}})$.
The starting point of our research was the following result, which deals with the quadratic number fields $K=\mathbb{Q}(\sqrt{d})$ such that $E_{\mathrm{tors}}(\mathbb{Q})\neq E_{\mathrm{tors}}(K)$, that is, the inclusion $E_{\mathrm{tors}}(\mathbb{Q})\subset E_{\mathrm{tors}}(K)$ is strict.
Conjecture 1.
This statement will be proved in the paper, in a case–by–case study, along with some other technical results (see Section 6).
In fact, let us assume $E_{\mathrm{tors}}(\mathbb{Q})\neq E_{\mathrm{tors}}(K)$ and consider $P\in E_{\mathrm{tors}}(K)\setminus E_{\mathrm{tors}}(\mathbb{Q})$ with order $n$. Then, from [6, 5], the primes which can divide the order of $P$ are $\{2,3,5,7\}$ and therefore if $\ell|n$, $\ell\in\{2,3,5,7\}$.
Since $P\in E[n]\cap E(K)$, we have that the coordinates of $P$ belong to $K\cap\mathbb{Q}(E[n])$. But $K\cap\mathbb{Q}(E[n])$ must be either $\mathbb{Q}$ or $K$, since $K/\mathbb{Q}$ is a quadratic extension. Since we are assuming that $P\notin E(\mathbb{Q})$, it follows that $K\subset\mathbb{Q}(E[n])$. On the other hand, we know from the Neron–Ogg–Shafarevich Criterion that $\mathbb{Q}(E[n])/\mathbb{Q}$ only ramifies at primes dividing $N_{E}$ or $n$ [11, Thm. 7.1]. Therefore, we can state:
@@
Let us take $K=\mathbb{Q}(\sqrt{d})$. Since every prime dividing $d$ ramifies in $K$, we have the following more appropriate version of the previous proposition.
@@
Let $E/\mathbb{Q}$ be an elliptic curve. If $K$ is a quadratic number field such that $E_{\mathrm{tors}}(\mathbb{Q})\not=E_{\mathrm{tors}}(K)$, then $K=\mathbb{Q}(\sqrt{d})$, where the primes dividing $d$ belong to the set $\{2,3,5,7\}\cup\{p:p|N_{E}\}$.
We have two possible situations concerning the growth of torsion. It may happen that a new torsion point $P\in E(K)$ appears with no nontrivial multiple of $P$ belonging to $E(\mathbb{Q})$ (the strict case), or it may happen that for all new torsion points $P\in E(K)$, there is a nontrivial multiple $mP$ belonging to $E(\mathbb{Q})$ (the mixed case).
@@
Note that, in the strict cases, we can always assume that the order of $P$ is a prime number $\ell$ since, if such a point $P$ appears, some multiple of it $mP$ will have prime order, and under our assumption it can not be contained in $E_{\mathrm{tors}}(\mathbb{Q})$.
@@
The mixed case can in principle involve points of any prime order in $\{2,3,5,7\}$ but, under our hypothesis, at least one new torsion point in this case must have an order that is a multiple of $2$, see Theorem 2 in [2].
Our take on the above conjecture will be as follows: Fix $E$ as in the conjecture above and assume $\ell\nmid N_{E}$. We will study, case by case, how the situation unfolds for the possible primes $\ell\in\{2,3,5,7\}$. In particular:
  1. ($\ell=2$)
    In this case we distinguish between the strict case and the mixed case. In both cases we perform a careful analysis of the valuation of the coordinates of the torsion point, together with some other relevant quantities.
  2. ($\ell=3$)
    This prime has a very specific behaviour as it is the only one who allows unexpected points to appear. Our way of approaching its description will be through the study of the Galois representation and a very detailed listing of available subgroups. As a consequence, more detailed results can be presented on how and where these points might be expected.
  3. ($\ell=5,7$)
@@
    These two primes can share a common approach via the study of the inertia group of a prime $\Lambda|\ell$ of $\mathbb{Q}(E[\ell])$. Essentially, the size of the groups involved will render impossible the existence of points of order $\ell$ in $E_{\mathrm{tors}}(K)$, although it will need to be proved in a case–by–case argument, depending on the reduction type of $E$ at $\ell$.
This paper is divided into six sections. Sections $2$ and $3$ are devoted to the prime $\ell=2$ (strict and mixed case, respectively), Section 4 is devoted to the prime $\ell=3$ (we only need to consider the strict case), and Section 5 is devoted to the primes $\ell=5,7$ (again, we only need to address the strict case). This section is independent from Sections 2, 3 and 4. The last section collects the results from the previous sections in order to state our main result. To facilitate the reading of the paper, we have included an appendix where we collect some information on subgroups of $\mathrm{GL}_{2}(\mathbb{F}_{5})$ and $\mathrm{GL}_{2}(\mathbb{F}_{7})$.
Examples are taken from [8] and labeled accordingly.
@@ 2. The prime $\ell=2$: The strict case
${\mathbb{Q}(E[2])\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\overline{\rho}_{E,2}(G_{\mathbb{Q}})}$$\scriptstyle{H}$${K\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{Q}}$
@@
$G_{4}=\mathrm{GL}_{2}(\mathbb{F}_{2})=\left\{\mathrm{Id},\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\begin{pmatrix}1&1\\ 1&0\end{pmatrix},\begin{pmatrix}0&1\\ 1&1\end{pmatrix},\begin{pmatrix}1&0\\ 1&1\end{pmatrix},\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\right\}$
@@
  • Curve $315.a$ 3: This curve has Galois group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[2])/\mathbb{Q})\simeq G_{2}$, and $E_{\mathrm{tors}}(\mathbb{Q})\simeq\mathcal{C}_{2}$. Over the field $\mathbb{Q}(\sqrt{-5})$, it obtains torsion $\mathcal{C}_{2}\times\mathcal{C}_{2}$.
  • Curve $317.a$ 3: This curve has Galois group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[2])/\mathbb{Q})\simeq G_{2}$, and $E_{\mathrm{tors}}(\mathbb{Q})\simeq\mathcal{C}_{4}$. Over the field $\mathbb{Q}(i)$, it obtains torsion $\mathcal{C}_{4}\times\mathcal{C}_{2}$.
  • Curve $815.a$ 8: This curve has Galois group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[2])/\mathbb{Q})\simeq G_{2}$, and $E_{\mathrm{tors}}(\mathbb{Q})\simeq\mathcal{C}_{8}$. Over the field $\mathbb{Q}(i)$, it obtains torsion $\mathcal{C}_{8}\times\mathcal{C}_{2}$.
@@
is an odd number. The expressions for the integers $b_{i}$ in terms of the integers $a_{i}$ can be found, for example, in [11, Ch. III].
@@
The $x$–coordinates of the nontrivial $2$-torsion points are the three roots of this polynomial, say $\alpha$, $\beta$, and $\gamma$, and we have $\mathbb{Q}(E[2])=\mathbb{Q}(\alpha,\beta,\gamma)$.
@@
Using again that the polynomial in Equation (2) lies in $\mathbb{Z}[X]$, we have $-4\alpha\beta c\in\mathbb{Z}$. Taking $2$-adic valuation, we obtain
(3) $2+2v+w\geq 0$
@@
Let us observe that
$v=v_{2}(a+b\sqrt{d})\geq\min\left\{v_{2}(a),v_{2}\big{(}b\sqrt{d}\big{)}\right\}.$
As $d$ is square-free we can assume from now on, with no loss of generality, that $v_{2}(d)=1$ Notice that in this case $v(a)\neq v(b\sqrt{d})$ and hence we have $v=\min\{v_{2}(a),v_{2}(b\sqrt{d})\}$; hence $v\leq v_{2}(a)$ and $v\leq v_{2}(b\sqrt{d})$.
Let us finally aim for a contradiction. We will distinguish three cases:
$w=-2$ In this case, Equation (3) implies that $v\geq 0$. Now, let us evaluate equation (4): we can see the expression
$a^{2}-b^{2}d-2ac+c^{2}$
as the sum of three terms:
  • $\circ$
    $a^{2}-b^{2}d$, which has valuation $2v\geq 0$,
  • $\circ$
    $-2ac$, which has valuation $11+v_{2}(a)+w\geq 1+v-2\geq-$ 1, and
  • $\circ$
    $c^{2}$, which has valuation $42w=-$ 4.
Therefore, the valuation of the term $a^{2}-b^{2}d-2ac+c^{2}$ is $-4$, and equation (4) results in
$0=6+v_{2}(b^{2}d)-8=v_{2}(b^{2}d)-2,$
but the valuation of $b^{2}d$ is always an odd number, hence we have a contradiction.
$w=-1$ In this case, Equation (3) implies that $v\geq-1/2$. We evaluate the equation (4): once again, we can see the expression
$a^{2}-b^{2}d-2ac+c^{2}$
as the sum of three terms:
  • $\circ$
    $a^{2}-b^{2}d$, which has a valuation of $12v\geq-$ 1,
  • $\circ$
    $-2ac$, which has a valuation of $21+v_{2}(a)+w\geq 1-(1/2)-1\geq-1/$ 2, and
  • $\circ$
    $c^{2}$, which has a valuation of $22w=-$ 2.
Therefore, the valuation of the term $a^{2}-b^{2}d-2ac+c^{2}$ is $-2$, and the equation (4) becomes
$0=6+v_{2}(b^{2}d)-4=v_{2}(b^{2}d)+2,$
but again the valuation of $b^{2}d$ is always an odd number, so we have a contradiction.
$w\geq 0$ In this case, Equation (3) implies that $v\geq-1$. Notice that $-1\leq v\leq v_{2}(b\sqrt{d})$; since the latter number is of the form $1/2+n$ with $n\in\mathbb{Z}$, we have $-1/2\leq v_{2}(b\sqrt{d})$.
Once again, we examine equation (4):
$a^{2}-b^{2}d-2ac+c^{2}$
is the sum of three terms:
  • $\circ$
    $a^{2}-b^{2}d$, which has a valuation of $22v\geq-$ 2;
  • $\circ$
    $-2ac$, which has a valuation of $1+v_{2}(a)+w\geq 1+v+0\geq 0$, and
  • $\circ$
    $c^{2}$, which has a valuation of $2w\geq 0$.
Therefore, the term $b^{2}d-2ac+c^{2}$ has a valuation greater than or equal to $-2$, and it results in
$0\geq 6+v(b^{2}d)-4=v(b^{2}d)+2\geq 2(-1/2)+2=1,$
which is a contradiction. This finishes the proof. ∎

### v2

@@ Abstract.
This work was supported by IMUS-Maria de Maeztu grant CEX2024-001517-M - Apoyo a Unidades de Excelencia María de Maeztu, funded by MICIU/AEI/ 10.13039/501100011033, grant PID2024-156912N funded by MICIU/AEI/10.13039/501100011033 and ERDF/EU and grants SOL2024-31596 and SOL2024-31708 from the Plan Propio de Investigación y Transferencia of the University of Sevilla, cofunded by the EU - Ministerio de Hacienda y Función Pública - Fondos Europeos - Junta de Andalucía – Consejería de Universidad, Investigación e Innovación”.
@@ 1. Introduction: The problem
  • We will write $o(P)$ for the order of the point $P$ on the group $E(K)$.
  • Whenever we consider a quadratic number field written as $K=\mathbb{Q}(\sqrt{d})$ we will assume $d$ is a square-free integer.
  • Examples are taken from [9] and labeled accordingly.
Let $E/\mathbb{Q}$ be an elliptic curve defined over $\mathbb{Q}$. Let $\ell$ be a prime number and let us write:
@@
This paper deals with properties of the torsion subgroup of elliptic curves defined over the rationals under quadratic field extensions. Let then $E/\mathbb{Q}$ be an elliptic curve, $K$ a quadratic number field. The first papers addressing the comparison between $E_{\mathrm{tors}}(\mathbb{Q})$ and $E_{\mathrm{tors}}(K)$ were [3, 4, 11]. In particular, [3, Theorem 2] states the following:
Theorem 1.
With the previous notations, we have the following table:
[…]
In the same reference, three further problems were stated. Problems 1 and 3 from this list were solved in [4, 11] independently. The origin of our research was trying to answer the problem originally stated in [3, Section 4] as Problem 2.
Problem: Is there a precise (and easy) description of which are the possible quadratic number fields $K=\mathbb{Q}(\sqrt{d})$ with $E_{\mathrm{tors}}(\mathbb{Q})\neq E_{\mathrm{tors}}(K)$, ideally in terms of some invariant(s) of the curve?
The extensive calculations from [4] were the starting point of our research, as we tried to prove conditions on the integer $d$ (alternatively, on the field $K$), assuming we do have growth of torsion subgroups. More specifically, we were interested on sieving the possible prime divisors of $d$ from invariants of $E$.
Let now $E/\mathbb{Q}$ be an elliptic curve with conductor $N_{E}$ and $K=\mathbb{Q}(\sqrt{d})$ a quadratic number field with $E_{\mathrm{tors}}(\mathbb{Q})\neq E_{\mathrm{tors}}(K)$. Assume $P\in E_{\mathrm{tors}}(K)\setminus E_{\mathrm{tors}}(\mathbb{Q})$ with $o(P)=n$. We know from [7, 6] the primes $\ell$ which can divide $n$ are $\{2,3,5,7\}$.
Since $P\in E[n]\cap E(K)$, we have that the coordinates of $P$ belong to $K\cap\mathbb{Q}(E[n])$. But $K\cap\mathbb{Q}(E[n])$ must be either $\mathbb{Q}$ or $K$, since $K/\mathbb{Q}$ is a quadratic extension. As we are assuming that $P\notin E(\mathbb{Q})$, it follows that $K\subset\mathbb{Q}(E[n])$. On the other hand, we know from the Neron–Ogg–Shafarevich Criterion that $\mathbb{Q}(E[n])/\mathbb{Q}$ only ramifies at primes dividing $N_{E}$ or $n$ [14, Theorem 7.1]. Therefore, we can state:
@@
Remark 1.
Under the hypothesis of the proposition, if $E$ has good reduction at $p\in\{2,3,5,7\}$ (i.e. $p\nmid N_{E}$), then $p$ ramifying in $K$ implies $E_{\mathrm{tors}}(\mathbb{Q})[p]\neq E_{\mathrm{tors}}(K)[p]$.
Let us now write $K=\mathbb{Q}(\sqrt{d})$. Since every prime dividing $d$ ramifies in $K$, we have the following version of the previous proposition, more appropriate for our goals.
@@
Let $E/\mathbb{Q}$ be an elliptic curve. If $K$ is a quadratic number field such that $E_{\mathrm{tors}}(\mathbb{Q})\not=E_{\mathrm{tors}}(K)$, then we can write $K=\mathbb{Q}(\sqrt{d})$, where the primes dividing $d$ belong to the set $\{2,3,5,7\}\cup\{p:p|N_{E}\}$.
A first property we looked into after this was the following natural sequel: If $p\in\mathbb{Z}$ is a prime such that $p|d$, can we always say $p|N_{E}$? As stated before, under these circumstances, either $p|N_{E}$ or $E_{\mathrm{tors}}(\mathbb{Q})[p]\neq E_{\mathrm{tors}}(K)[p]$, so it is natural to look at points defined over $K$ of prime order. Our strategy of choice for that study has been the so–called mod $\ell$ Galois representations, which we review now briefly.
Let $\ell$ be a prime integer. As it is well–known [13], the action of the absolute Galois group $\mathrm{G}_{\mathbb{Q}}=\mathop{\mathrm{Gal}}\nolimits(\overline{\mathbb{Q}}/\mathbb{Q})$ on $E[\ell]$ defines a mod $\ell$ Galois representation
The extension $\mathbb{Q}(E[\ell])/\mathbb{Q}$ is Galois, with $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[\ell])/\mathbb{Q})\simeq\overline{\rho}_{E,\ell}(\mathrm{G}_{\mathbb{Q}})$. In this context, therefore, the subgroup lattice of $\mathop{\mathrm{GL}}\nolimits_{2}(\mathbb{F}_{\ell})$ offers relevant information about the field extensions we are interested in.
Fix then $E/\mathbb{Q}$ an elliptic curve, $K=\mathbb{Q}(\sqrt{d})$ a quadratic number field such that $E_{\mathrm{tors}}(\mathbb{Q})\not=E_{\mathrm{tors}}(K)$ and $\ell$ a prime such that $\ell\nmid N_{E}$.
We will study, case by case, how the situation unfolds for the possible primes $\ell\in\{2,3,5,7\}$ separatedly.
The case $\ell=2$. (Sections 2 & 3). In this case we were able to prove the following.
Theorem 2.
Let $E/\mathbb{Q}$ be an elliptic curve, and $K$ a quadratic number field such that $E_{\mathrm{tors}}(\mathbb{Q})\not=E_{\mathrm{tors}}(K)$. If $2|d$, then $2|N_{E}$.
The proof of this theorem encompasses Propositions 3, 4, 5 and 6, which also yield some collateral results as Corollary 1. It is the most complicated result, and it must be broken, to begin with, in two parts, illustrating two different situations for the torsion growth.
We know a new torsion point must appear when moving from $\mathbb{Q}$ to $K$. It may happen that some new torsion point $P\in E(K)$ appears with no nontrivial multiple of $P$ belonging to $E(\mathbb{Q})$ (this will be called the strict case). Or it may happen that, for each new torsion point $P\in E(K)$, there is a nontrivial multiple $mP$ belonging to $E(\mathbb{Q})$ (this will be the mixed case).
@@
Note that, in the strict cases, we can always assume that the order of $P$ is a prime number $\ell$ since, if such a point $P$ appears, some multiple of it will have prime order, and under our assumption it can not be contained in $E_{\mathrm{tors}}(\mathbb{Q})$.
@@
The mixed case will be the most involved and it is distinctive of the $\ell=2$ case. Indeed, this situation can in principle happen with points of any prime order in $\{2,3,5,7\}$ but, under our hypothesis, at least one new torsion point in this case must have an order that is a multiple of $2$, as shown in Theorem 1.
The case $\ell=3$. (Section 4). This prime has a very specific behaviour as it is the only one who allows unexpected points to appear. By that we mean we know for sure that $3|d$ does not imply $3|N_{E}$. An example of this is the curve 19.a2, which verifies
$N_{E}=19,\qquad E_{\mathrm{tors}}(\mathbb{Q})=\mathcal{C}_{1},\qquad E_{\mathrm{tors}}\big(\mathbb{Q}(\sqrt{-3})\big)=\mathcal{C}_{3}.$
Therefore what we aim for here is a more complete understanding of the phenomenon, and give conditions under which we can be sure that $3|d$ implies $3|N_{E}$. This is still a work in progress but we have been able to prove results on this regard, like the following ones (appearing in Section 4 as Propositions 7 and 8):
Proposition 7. Let $E/\mathbb{Q}$ be an elliptic curve such that $E_{\mathrm{tors}}(\mathbb{Q})$ is trivial and $E_{\mathrm{tors}}(K)=\mathcal{C}_{9}$, with $K$ a quadratic field. Then, $E$ has bad reduction at 3.
Proposition 8. Let $E/\mathbb{Q}$ be an elliptic curve and $K=\mathbb{Q}(\sqrt{d})$ be a quadratic extension such that there exists a point $P\in E[3]$ such that $P\in E(K)\setminus E(\mathbb{Q})$. Let us assume $3|d$ and $E$ has good reduction at 3 ($3\nmid N_{E}$), then
$E_{\mathrm{tors}}(K)=E_{\mathrm{tors}}(\mathbb{Q})\times\mathcal{C}_{3}.$
The cases $\ell=5,7$. (Section 5). As in the $\ell=2$ case we have here:
Theorem 3.
Let $E/\mathbb{Q}$ be an elliptic curve, $K$ a quadratic number field such that $E_{\mathrm{tors}}(\mathbb{Q})\neq E_{\mathrm{tors}}(K)$. If $p=5,7$ ramifies in $K$, then $p|N_{E}$.
@@
This result was already known and it can be proved from [12, Theorem 1] (using dd–twists) and also from [10, Theorem 1.5] (using reduction types in abelian varieties over more general base fields). Nevertheless we have applied the mod $\ell$ Galois representation techniques in order to (hopefully) find a new proof which would allow us to better understand the phenomenon. In particular, these two primes can share a common approach via the study of the inertia group of a prime $\Lambda|\ell$ of $\mathbb{Q}(E[\ell])$. Essentially, the size of the groups involved will render impossible the existence of points of order $\ell$ in $E_{\mathrm{tors}}(K)$, although it will need to be proved in a case–by–case argument, depending on the reduction type of $E$ at $\ell$.
Besides succeeding in finding an alternative proof, our arguments can be used to prove the following result, which cannot be derived from [12, 10] and is, in fact, a stronger statement than Theorem 3:
Theorem 4.
Let $E/\mathbb{Q}$ be an elliptic curve. Assume $K$ is a quadratic number field such that there exists $P\in E(K)[\ell]\setminus E(\mathbb{Q})$, with $\ell\geq 3$ prime.
  1. (1)
    For every prime $p\neq\ell$ that ramifies in $K$, $E$ has additive reduction at $p$, i.e. $p^{2}|N_{E}$.
  2. (2)
    If $p=\ell>3$ ramifies in $K$, $E$ has additive reduction at $p$, i.e. $p^{2}|N_{E}$.
Section 6, finally, is devoted to conclusions and final remarks, including comments on future directions of our work. To make the reading of the paper a little easier, we have included an appendix where we collect some information on subgroups of $\mathrm{GL}_{2}(\mathbb{F}_{5})$ and $\mathrm{GL}_{2}(\mathbb{F}_{7})$ which are needed in Section 5.
The authors want to thank the anonymous referees for their insights which actually allowed us to realise Theorem 4 and which have meant an important contribution to the final form of this paper, often by shortening arguments. They also want to thank E. González–Jiménez and F. Najman for their conversations on the topic and support during the preparation of the paper. Finally, thanks are also due to M. Melistas who pointed us out his contributions on the matter.
@@ 2. The prime $\ell=2$: The strict case
[…]
@@
$G_{4}=\mathrm{GL}_{2}(\mathbb{F}_{2})=\left\{\mathrm{Id},\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\begin{pmatrix}1&1\\ 1&0\end{pmatrix},\begin{pmatrix}0&1\\ 1&1\end{pmatrix},\begin{pmatrix}1&0\\ 1&1\end{pmatrix},\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\right\}$.
@@
  • Curve $15.a3$: This curve has Galois group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[2])/\mathbb{Q})\simeq G_{2}$, and $E_{\mathrm{tors}}(\mathbb{Q})\simeq\mathcal{C}_{2}$. Over the field $\mathbb{Q}(\sqrt{-5})$, it obtains torsion $\mathcal{C}_{2}\times\mathcal{C}_{2}$.
  • Curve $17.a3$: This curve has Galois group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[2])/\mathbb{Q})\simeq G_{2}$, and $E_{\mathrm{tors}}(\mathbb{Q})\simeq\mathcal{C}_{4}$. Over the field $\mathbb{Q}(i)$, it obtains torsion $\mathcal{C}_{2}\times\mathcal{C}_{4}$.
  • Curve $15.a8$: This curve has Galois group $\mathop{\mathrm{Gal}}\nolimits(\mathbb{Q}(E[2])/\mathbb{Q})\simeq G_{2}$, and $E_{\mathrm{tors}}(\mathbb{Q})\simeq\mathcal{C}_{8}$. Over the field $\mathbb{Q}(i)$, it obtains torsion $\mathcal{C}_{2}\times\mathcal{C}_{8}$.
@@
is an odd number. The expressions for the integers $b_{i}$ in terms of the integers $a_{i}$ can be found, for example, in [14, Chapter III].
@@
The xx–coordinates of the nontrivial $2$-torsion points are the three roots of this polynomial, say $\alpha$, $\beta$, and $\gamma$, and we have $\mathbb{Q}(E[2])=\mathbb{Q}(\alpha,\beta,\gamma)$.
@@
Using again that the polynomial in Equation (2) lies in $\mathbb{Z}[X]$, we have $4\alpha\beta c\in\mathbb{Z}$. Taking $2$-adic valuation, we obtain
(3) $2+2v+w\geq 0.$
@@
As $b\in\mathbb{Q}$, this equation implies that $1=-v_{2}(d)=6+2v_{2}(b)+2v_{2}(a^{2}-b^{2}d-2ac+c^{2})$ is even, a contradiction. ∎